r2 - 10 May 2011 - 13:26:29 - KaushikDeYou are here: TWiki >  Admins Web > MinutesDataManageMay10



Minutes of the US ATLAS Data Management meeting, May 10, 2011
  • Previous meetings and background : IntegrationProgram
  • Coordinates: Tuesdays, Noon Central
    • (570) 310-0130, Access code: 735188; Dial *6 to mute/un-mute.


  • Meeting attendees: Saul, Armen, Dave, Charles, Wei, Michael, Shawn, Patrick, Wensheng, John
  • Apologies:
  • Guests:

Topics for this week

  • Storage status - Armen
    • Everything looks ok.
    • Illinois PRODDISK close to full. Dave - waiting for LFC dump from Hiro. UC may need to run pandamover cleanup.
    • MWT2 LOCALGROUPDISK has been cleaned up - thanks.
    • SWT2 GROUPDISK 133TB used - more than twice other sites?
  • Central deletion - Hiro, Armen, Charles
    • Rate has been low at BNL, Michael does not know of any reason at BNL - Armen will follow up with ADC.
    • Need to cleanup dark data on MCDISK.
  • Storage reporting - Dave, Hiro
    • Test unallocated, unpowered tokens.
  • Hot issues
    • Wei - Bestman 2 is working much better in recovering from storage errors. No problem so far like used to happen with Bestman 1.
  • AOB
    • Patrick - CPB coming next week.
    • Hiro - deployment of new SS at BNL.
    • Tier 3 consistency reports - who and how fixes them?

-- KaushikDe - 10 May 2011

About This Site

Please note that this site is a content mirror of the BNL US ATLAS TWiki. To edit the content of this page, click the Edit this page button at the top of the page and log in with your US ATLAS computing account name and password.


Powered by TWiki
This site is powered by the TWiki collaboration platformCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding TWiki? Send feedback